# RMMI: Enhanced Obstacle Avoidance for Reactive Mobile Manipulation using an Implicit Neural Map

### v1

@@ Abstract
We introduce RMMI, a novel reactive control framework for mobile manipulators operating in complex, static environments. Our approach leverages a neural Signed Distance Field (SDF) to model intricate environment details and incorporates this representation as inequality constraints within a Quadratic Program (QP) to coordinate robot joint and base motion. A key contribution is the introduction of an active collision avoidance cost term that maximises the total robot distance to obstacles during the motion. We first evaluate our approach in a simulated reaching task, outperforming previous methods that rely on representing both the robot and the scene as a set of primitive geometries. Compared with the baseline, we improved the task success rate by 25% in total, which includes increases of 10% by using the active collision cost. We also demonstrate our approach on a real-world platform, showing its effectiveness in reaching target poses in cluttered and confined spaces using environment models built directly from sensor data. For additional details and experiment videos, visit https://rmmi.github.io/.
Index Terms:
Mobile Manipulation, Collision Avoidance, Motion Control and Reactive Control
@@ I Introduction
Traditionally, reactive control has been valued for its speed and adaptability, though it was often considered less capable in handling complex environments compared to motion planning, which, while more capable, is typically slower and computationally intensive. Recent works in motion planning [3, 4] have made significant strides in improving computational efficiency, allowing for faster planning in dynamic and cluttered environments. These advancements are narrowing the gap between reactive control and motion planning, making them more comparable in terms of speed and capability.
Figure 1: Illustration of our approach to reactive control in cluttered environments. The obstacles are represented within a Signed Distance Field (SDF), shown in red, where darker shades indicate closer proximity to the obstacle surfaces. The robot is represented as a set of points sampled from its outer geometry (depicted as magenta dots). Our controller computes the joint velocities required to follow a trajectory from the start to the end pose (highlighted in yellow), while simultaneously maximising manipulability and maintaining a safe distance from obstacles, all within the constraints of joint limits and collision avoidance.
@@
To address these limitations, we propose a novel reactive controller that incorporates collision avoidance by querying a 3D representation of the robot’s surroundings, modelled as a neural Signed Distance Field (SDF). Unlike traditional methods, our approach leverages the continuous representation provided by the neural SDF, allowing for precise obstacle detection and efficient collision avoidance in cluttered environments. Another advantage of using neural SDFs are the fixed query times that lead to a consistent control rate.
Our method enhances collision avoidance by representing the robot’s outer geometry with sampled points. These points are queried against the SDF to determine their proximity to obstacles, and the resulting distance and direction data are incorporated into the controller’s optimisation as inequality constraints. Additionally, we introduce an active collision avoidance cost within the optimisation framework, which further increases the robot’s safety margin by maximising its distance from obstacles during motion. Notably, as shown in Fig. 1, our controller enables the robot to perform complex tasks, such as reaching target poses inside a bookshelf, with enhanced precision.
In summary, the contributions of this letter are:
  1. 1.
    The formulation of a reactive controller that incorporates collision avoidance in 3D, in which the robot is represented as a set of points sampled from its outer mesh. By obtaining the distance of each point using a neural Signed Distance Field, we formulate different inequality constraints that are incorporated into the controller optimisation problem.
  2. 2.
    The introduction of an active collision avoidance cost to the optimisation problem that steers the robot away from obstacles by averaging the individual distances projected in the robot’s configuration space. The active collision avoidance cost leads to a further average performance increase of 10% in task completion.
  3. 3.
    A thorough validation of our approach through simulated and real-world experiments, demonstrating its effectiveness in performing precise mobile manipulation tasks such as reaching target poses inside a bookshelf and behind objects on a cluttered table [^1] Videos and 3D reconstructions are available on https://rmmi.github.io/. We also ablate the various components within our framework.
@@ II-A Collision Avoidance in Motion Generation
While effective, these methods often rely on simplified collision primitives or precise object models, limiting their applicability in unstructured environments [11, 5]. To address these limitations, some researchers have explored sensor-based approaches, using data from 2D laser scans on mobile manipulators [7, 6] or depth maps for static manipulators [13, 12]. However, these methods face challenges: 2D laser scans restrict the space in which a robot arm can operate, while depth maps are less practical for mobile manipulators due to limited viewing angles from robot-mounted cameras.
Our work advances these approaches by integrating a 3D representation that efficiently models collisions within a holistic reactive controller, eliminating the need for predefined geometric models and overcoming some of the limitations of purely sensor-based methods. The next section delves into the challenge of developing efficient representations for collision avoidance within the context of motion generation.
@@ II-C Signed Distance Fields for Collision Checking
Moreover, some researchers have explored modelling the robot itself as an SDF by learning the robot’s SDF using neural networks [27] or basis functions [28]. The robot’s SDF enables precise control by computing the distance between the robot and its environment using static RGB-D cameras [13] or external tracking systems [28, 27].
@@ III-A Holistic Mobile Manipulator Motion Controller
At the core of this controller is a Quadratic Program (QP) that is designed to output the required joint velocities ${\dot{{\mathitbf{q}}}}\in\mathbb{R}^{n}$, where $n$ represents the degrees of freedom of the mobile manipulator, including both the arm and the base. The primary objective is to achieve the desired spatial velocity ${\mathitbf{v}}\in\mathbb{R}^{6}$ of the end-effector in the base frame ${}^{b}{\mathitbf{v}}_{e}^{*}$, where $e$ represents the end-effector frame and $b$ the robot’s base frame.
The optimisation problem is defined as the following QP formulation:
@@
[…]
@@
with ${{\bm{J}}}_{m}\in\mathbb{R}^{n}$ representing the manipulability Jacobian [29] that maximises the manipulability index of the robot arm [30], which indicates how well conditioned the current arm configuration is. The term ${{\bm{J}}}_{o}\in\mathbb{R}^{n}$ compromises a linear component to optimise the base’s relative orientation to the end-effector [1].
@@
[…]
where $\,{}^{\scriptscriptstyle b}\!\,{{\bm{J}}}_{e}\in\mathbb{R}^{6\times n}$ maps joint velocities to the end-effector’s spatial velocity.
@@ III-B Signed Distance Fields
Signed distance fields (SDF) are a well-studied representation of objects and environments that is widely used in robotics [17], computer vision [31], and computer graphics [32]. In the 3D case, an SDF is a function $f:\mathbb{R}^{3}\rightarrow\mathbb{R}$ that returns the distance $d\in\mathbb{R}$ of a sample point ${\mathitbf{p}}\in\mathbb{R}^{3}$ to the nearest surface. The sign of this distance indicates if the sampled point ${\mathitbf{p}}$ is inside the surface (negative), outside (positive), or on the surface (zero distance). Consequently, the surface $\mathcal{S}$ of the obstacle is represented by the zero-level set of the SDF:
$\mathcal{S}=\left\{{\mathitbf{p}}\in\mathbb{R}^{3}|f({\mathitbf{p}})=0\right\}.$ (8)
In addition, we can determine the direction to the closest obstacle by computing the gradient ${\hat{{\mathitbf{g}}}}\in\mathbb{R}^{3}$ of this function. This gradient vector points from the sampling point away from the surface, enabling the closest point on the surface ${\mathitbf{p}}_{\text{near}}$ to be estimated as [33]:
@@ IV-A Robot Representation
This operation results in a set of $N$ points, where $N=\sum_{k}M_{k}$ represents the total number of points sampled from the entire robot. An illustration comparing this points-based representation to a traditional sphere-based approach [4] is shown in Fig. 2, highlighting the precision of our method in capturing the robot’s geometry.
@@ IV-B Collision Constraints
Here, $\,{}^{\scriptscriptstyle w}\!{\dot{{\mathitbf{p}}}}(t)$ represents the translational velocity of the point, which can be expressed using the manipulator Jacobian ${{\bm{J}}}_{\nu}(\tilde{{\mathitbf{q}}})\in\mathbb{R}^{(3\times k)}$ considering each point as the end-effector [29], where $\tilde{{\mathitbf{q}}}=\left(q_{0},q_{1},...q_{k}\right)\in\mathbb{R}^{k}$ denotes the joints up to the $k$-th link. Rewriting $\,{}^{\scriptscriptstyle w}\!{\dot{{\mathitbf{p}}}}(t)$ with respect to the joint’s velocity results in:
@@
${\hat{{\mathitbf{g}}}}\cdot{{\bm{J}}}_{\nu}(\tilde{{\mathitbf{q}}})\cdot{\tilde{{\mathitbf{q}}}}=\dot{d}(t),$ (15)
@@
${{\bm{J}}}_{d_{j}}(\tilde{{\mathitbf{q}}_{j}})={\hat{{\mathitbf{g}}}}_{j}\cdot{{\bm{J}}}_{\nu_{j}}(\tilde{{\mathitbf{q}}_{j}}).$ (16)
Using a general velocity damper [34], we impose the following inequality:
@@
where $d_{i}$ is the influence distance, i.e. the distance where the point is considered in the optimisation, and $d_{s}$ is the stopping distance, i.e. the minimum allowable distance to an obstacle.
@@
$\begin{pmatrix}{{\bm{J}}}_{d_{1}}(\tilde{{\mathitbf{q}}_{1}})\tilde{{\mathitbf{q_{1}}}}&\mathbf{0}_{1\times 6+n-k_{1}}\\ \vdots&\vdots\\ {{\bm{J}}}_{d_{N}}(\tilde{{\mathitbf{q}}_{N}})\tilde{{\mathitbf{q}}_{N}}&\mathbf{0}_{1\times 6+n-k_{N}}\\ \end{pmatrix}{\mathitbf{x}}(t)\leq\begin{pmatrix}\frac{f({\mathitbf{p}}_{1})-d_{s}}{d_{i}-d_{s}}\\ \vdots\\ \frac{f({\mathitbf{p}}_{N})-d_{s}}{d_{i}-d_{s}}\end{pmatrix}$ (18)
where the matrix on the left side is defined as ${{\bm{A}}}_{c}\in\mathbb{R}^{N\times n+6}$, and the right side vector as ${\mathitbf{b}}_{c}\in\mathbb{R}^{N}$. Both can be integrated into the QP via Eq. 3.
@@ IV-C Active Collision Avoidance
To formulate this cost, we define a collision Jacobian ${{\bm{J}}}_{c}$ to represent how the distance between the robot and obstacles changes with respect to the joint velocities. This Jacobian is calculated as a weighted average of the individual distance Jacobians ${{\bm{J}}}_{d_{j}}$, with weights proportional to the distances:
@@
However, ${{\bm{J}}}_{c}$ is a normalised direction; to consider a cost that is dependent on the robot distance, we introduce a dynamic gain $\lambda_{c}$, inspired by the collision cost in CHOMP [18]:
$\lambda_{c}(t)=\frac{\lambda^{max}_{c}}{\left(d_{i}-d_{s}\right)^{2}}\cdot\left(r_{d}(t)-d_{i}\right)^{2},$ (21)
where $r_{d}$ is the distance of the robot as $\min_{j}f(\,{}^{\scriptscriptstyle w}\!{\mathitbf{p}}_{j})$, and $\lambda^{max}_{c}$ controls the steepness of the active collision gain.
@@
This overall optimisation encourages the robot to maintain a safer distance from obstacles while optimising for manipulability and orientation.
@@ V-A2 Tasks
The first set of experiments involved a reaching task, simulated using the Swift simulator and the Robotics Toolbox for Python [35]. The SDFs representing the obstacles were generated by combining analytic distance functions of primitive geometries [19].
@@
We generated 500 different variations for each scenario, ensuring that any combination where the target pose resulted in a collision between the end-effector and obstacles was excluded.
@@ V-A3 Metrics
  • Success: The success rate of the controller in generating a collision-free trajectory that reaches the desired 6-DoF pose with a margin of 2 cm as previous work [36].
@@
  • Gracefullness: A proxy for the smoothness of motion, defined as being safe, comfortable, fast, and intuitive [37]. Following previous work [2], we measure the average absolute end-effector accelerations across different trials and environments, considering only successful motions to exclude cases where the robot failed due to local minima.
@@ V-A4 Baselines
We compare our approach against the holistic mobile manipulation controller presented in [1] and a modified version incorporating the collision avoidance component of NEO [5]. This provided a baseline for comparing the use of an implicit SDF versus mesh-to-mesh collision queries using primitives. Additionally, we tested our controller with two sets of sampled points and a modified version [^3] The modification considers adding the Omron LD-60 Base. of the manually placed spheres from [4] to represent the robot. Finally, because our active collision avoidance is built by combining multiple inequality constraints, we included an experiment using the baseline [1] that incorporates our active collision cost.

### v2

# RMMI: Reactive Mobile Manipulation using an Implicit Neural Map

@@ Abstract
Mobile manipulator robots operating in complex domestic and industrial environments must effectively coordinate their base and arm motions while avoiding obstacles. While current reactive control methods gracefully achieve this coordination, they rely on simplified and idealised geometric representations of the environment to avoid collisions. This limits their performance in cluttered environments. To address this problem, we introduce RMMI, a reactive control framework that leverages the ability of neural Signed Distance Fields (SDFs) to provide a continuous and differentiable representation of the environment’s geometry. RMMI formulates a quadratic program that optimises jointly for robot base and arm motion, maximises the manipulability, and avoids collisions through a set of inequality constraints. These constraints are constructed by querying the SDF for the distance and direction to the closest obstacle for a large number of sampling points on the robot. We evaluate RMMI both in simulation and in a set of real-world experiments. For reaching in cluttered environments, we observe a 25% increase in success rate. For additional details, code, and experiment videos, please visit https://rmmi.github.io/.
@@ I Introduction
Traditionally, reactive control has been valued for its speed and adaptability, though it was often considered less capable in handling complex environments compared to motion planning [3], which, while more capable, is typically slower and computationally intensive. Recent works in motion planning [4, 5] have made significant strides in improving computational efficiency, allowing for faster planning in dynamic and cluttered environments. These advancements are narrowing the gap between reactive control and motion planning, making them more comparable in terms of speed and capability.
Figure 1: Illustration of our approach to reactive control in cluttered environments. The obstacles are represented within a Signed Distance Field (SDF), shown in red, where darker shades indicate closer proximity to the obstacle surfaces. The robot is represented as a set of points sampled from its outer geometry (depicted as magenta dots). Our controller computes the joint velocities required to follow a trajectory from the start to the end pose (highlighted in yellow), while simultaneously maximising manipulability and maintaining a safe distance from obstacles, all while staying within the constraints of joint limits and collision avoidance.
@@
To address these limitations, we propose integrating a neural Signed Distance Field (SDF) as the underlying scene representation. Unlike traditional SDFs, neural SDFs provide a continuous representation of the environment, enabling more precise and efficient collision avoidance in cluttered settings.
The main contribution of this paper is RMMI, a reactive control framework for mobile manipulators that incorporates information from a neural SDF into the constraints of a Quadratic Program (QP) to enhance performance in complex environments. Additionally, we introduce an active collision avoidance cost, which encourages the robot to maintain greater distances from obstacles, enhancing safety and further improving performance. Through both simulation and real-world experiments, we demonstrate that our approach improves task success rates by at least 25% over prior reactive controllers [6, 1]. Furthermore, it enables effective motion in cluttered and confined spaces using environment models constructed directly from onboard sensor data.
@@ II-A Collision Avoidance in Motion Generation
While effective, these approaches often rely on simplified collision primitives or precise object models, limiting their applicability in unstructured environments [12, 6]. To address these limitations, some researchers have explored sensor-based approaches, using data from 2D laser scans on mobile manipulators [8, 7] or depth maps for static manipulators [14, 13]. However, these approaches face challenges: 2D laser scans restrict the space in which a robot arm can operate, while depth maps are less practical for mobile manipulators due to limited viewing angles from robot-mounted cameras.
Our work advances these approaches by integrating a 3D representation that efficiently models collisions within a holistic reactive controller, eliminating the need for predefined geometric models and overcoming some of the limitations of purely sensor-based methods.
@@ II-C Signed Distance Fields for Collision Checking
Complementing environment-based approaches, some researchers have explored representing the robot itself as an SDF by learning its geometry using neural networks [27] or basis functions [28]. The robot’s SDF enables precise control by computing the distance between the robot and its environment using static RGB-D cameras [14] or external tracking systems [28, 27].
@@ III-A Holistic Mobile Manipulator Motion Controller
At the core of this controller is a Quadratic Program (QP) that is designed to output the required joint velocities ${\dot{{\mathitbf{q}}}}\in\mathbb{R}^{n}$, where $n$ represents the degrees of freedom of the mobile manipulator, considering the robot base as one revolute followed by a one prismatic virtual joint. The primary objective is to achieve a desired spatial velocity ${\mathitbf{v}}\in\mathbb{R}^{6}$ of the end-effector in the base frame $^{b}{\mathitbf{v}}_{e}^{*}$, where $e$ and $b$ represent the end-effector and base frame. The optimisation problem is formulated as a QP as follows:
@@
$\text{s.t}\text{ }\mathcal{J}{\mathitbf{x}}$ $=\,{}^{\scriptscriptstyle b}\!{\mathitbf{v}}_{e}$ (2)
@@
with ${{\bm{J}}}_{m}\in\mathbb{R}^{n}$ representing the manipulability Jacobian [29] that maximises the manipulability index of the robot arm [30], which indicates how well conditioned the current arm configuration is. The term ${{\bm{J}}}_{o}\in\mathbb{R}^{n}$ comprises a linear component to optimise the base’s relative orientation to the end-effector [1].
@@
$\mathcal{J}=\begin{pmatrix}\,{}^{\scriptscriptstyle b}\!\,{{\bm{J}}}_{e}&{{\bm{1}}}_{6\times 6}\end{pmatrix}\in\mathbb{R}^{6\times(n+6)},$ (7)
where $^{\scriptscriptstyle b}\!\,{{\bm{J}}}_{e}\in\mathbb{R}^{6\times n}$ maps joint velocities to the end-effector’s spatial velocity relative to the base frame.
@@ III-B Signed Distance Fields
Signed distance fields (SDF) are a well-studied representation of objects and environments that is widely used in robotics [18], computer vision [31], and computer graphics [32]. In the 3D case, an SDF is a function $f\colon\mathbb{R}^{3}\rightarrow\mathbb{R}$ that returns the distance $d\in\mathbb{R}$ of a sample point ${\mathitbf{p}}\in\mathbb{R}^{3}$ to the nearest surface. The sign of this distance indicates if the sampled point $\mathitbf{p}$} is inside the surface (negative), outside (positive), or on the surface (zero distance). Consequently, the surface $\mathcal{S}$ of the obstacle is represented by the zero-level set of the SDF:
$\mathcal{S}=\left\{{\mathitbf{p}}\in\mathbb{R}^{3}\mid f({\mathitbf{p}})=0\right\}.$ (8)
Additionally, we can determine the direction to the closest obstacle by computing the gradient ${\hat{{\mathitbf{g}}}}\in\mathbb{R}^{3}$ of this function. This gradient vector points from the sampling point away from the surface, enabling the closest point on the surface ${\mathitbf{p}}_{\text{near}}$ to be estimated as [33]:
@@ IV-A Robot Representation
This operation results in a set of $N$ points, where $N=\sum_{k}M_{k}$ represents the total number of points sampled from the entire robot. An illustration comparing this point-based representation to a traditional sphere-based approach [5] is shown in Fig. 2, highlighting the precision of our method in capturing the robot’s geometry.
Selecting an appropriate value for $N$ involves balancing geometric fidelity and computational efficiency. With a smaller $N$, each sampled point must be expanded by a margin to approximate the robot’s full geometry, which limits how closely it can approach obstacles. In contrast, increasing $N$ enhances geometric accuracy but raises computational costs. Our method scales linearly with the number of sampled points; however, this overhead is mitigated by parallelized batch queries for distance computations.
@@ IV-B Collision Constraints
Here, $^{\scriptscriptstyle w}\!{\dot{{\mathitbf{p}}}}(t)$ represents the translational velocity of the point, which can be expressed using the manipulator Jacobian ${{\bm{J}}}_{\nu}(\tilde{{\mathitbf{q}}})\in\mathbb{R}^{3\times k}$ considering each point as the end-effector [29], where $\tilde{{\mathitbf{q}}}=\left(q_{0},q_{1},...q_{k}\right)\in\mathbb{R}^{k}$ denotes the joints up to the $k$-th link. Rewriting $^{\scriptscriptstyle w}\!{\dot{{\mathitbf{p}}}}(t)$ with respect to the joint’s velocity results in:
@@
${\hat{{\mathitbf{g}}}}\cdot{{\bm{J}}}_{\nu}(\tilde{{\mathitbf{q}}})\cdot\dot{\tilde{{\mathitbf{q}}}}=\dot{d}(t),$ (15)
@@
[…]
Where the subscript $j$ is now reintroduced for clarity. Using a general velocity damper approach [34], we impose the following inequality:
@@
where $d_{i}$ is the influence distance, i.e. the distance at which the point is considered in the optimisation, and $d_{s}$ is the stopping distance, i.e. the minimum allowable distance to an obstacle.
@@
$\begin{pmatrix}{{\bm{J}}}_{d_{1}}(\tilde{{\mathitbf{q}}}_{1})&\mathbf{0}_{1\times 6+n-k_{1}}\\ \vdots&\vdots\\ {{\bm{J}}}_{d_{N}}(\tilde{{\mathitbf{q}}}_{N})&\mathbf{0}_{1\times 6+n-k_{N}}\\ \end{pmatrix}{\mathitbf{x}}(t)\leq\begin{pmatrix}\frac{f({\mathitbf{p}}_{1})-d_{s}}{d_{i}-d_{s}}\\ \vdots\\ \frac{f({\mathitbf{p}}_{N})-d_{s}}{d_{i}-d_{s}}\end{pmatrix}$ (18)
where the matrix on the left side is defined as ${{\bm{A}}}_{c}\in\mathbb{R}^{N\times(n+6)}$, and the right side vector as ${\mathitbf{b}}_{c}\in\mathbb{R}^{N}$. Both can be integrated into the QP via Eq. 3.
@@ IV-C Active Collision Avoidance
To formulate this cost, we define a collision Jacobian ${{\bm{J}}}_{c}\in\mathbb{R}^{n}$ to represent how the distance between the robot and obstacles changes with respect to the joint velocities. This Jacobian is calculated as a weighted average of the individual distance Jacobians ${{\bm{J}}}_{d_{j}}$, with weights proportional to the distances:
@@
However, ${{\bm{J}}}_{c}$ represents the rate of change of distance to the obstacle with respect to joint velocities. To incorporate a cost that depends on the robot’s distance to the obstacle, we design a dynamic gain $\lambda_{c}$, inspired by the collision cost in CHOMP [3]:
$\lambda_{c}(t)=\frac{\lambda^{\max}_{c}}{\left(d_{i}-d_{s}\right)^{2}}\cdot\left(r_{d}(t)-d_{i}\right)^{2},$ (21)
where $r_{d}$ is the distance of the robot defined as $\min_{j}f(\,{}^{\scriptscriptstyle w}\!{\mathitbf{p}}_{j})$, and $\lambda^{\max}_{c}$ controls the steepness of the active collision gain.
@@
This modified optimisation encourages the robot to maintain a safer distance from obstacles while optimising for manipulability and orientation.
@@ V-A2 Tasks
The first set of experiments involved a reaching task, simulated using the Swift simulator and the Robotics Toolbox for Python [35] with a fixed step time of 0.05s (i.e, a control loop of 20Hz). The SDFs representing the obstacles were generated by combining analytic distance functions of primitive geometries [19]. Using the ground truth geometry allows us to isolate the evaluation of our method from perception and SDF-generation errors, which are assessed in our real-world experiments in Section VI.
@@
We generated 500 different variations for each scenario, excluding any combinations where the target pose would result in a collision between the end-effector and obstacles.
Our proposed benchmark is available at https://github.com/nmarticorena/frankie_planner, which includes an RRT planner to verify the feasibility of the scenes.
@@ V-A3 Metrics
  • Success: The success rate of the controller in generating a collision-free trajectory that reaches the desired 6-DoF pose with a margin of 2cm as in previous work [36].
@@
  • Gracefulness: Following previous work [2] we measure the average absolute end-effector acceleration as a proxy for the smoothness of motion [37]. We only consider successful motions to exclude cases where the robot stalled due to local minima.
@@ V-A4 Baselines
We compare our approach against the holistic mobile manipulation controller presented in [1] and a modified version incorporating the collision avoidance component of NEO [6]. This provided a baseline for comparing the use of an implicit SDF versus mesh-to-mesh collision queries using primitives. Additionally, we tested our controller with two sets of sampled points and a modified version of the manually placed spheres approach from [5] (modified to include the Omron LD-60 Base) to represent the robot. Finally, because our active collision avoidance is built by combining multiple inequality constraints, we included an experiment using the baseline [1] that incorporates our active collision cost.
V-A5 Hardware specification
All experiments were conducted on a machine with an Intel i7-12700K CPU, an NVIDIA RTX 3090 GPU, and 32GB RAM.